Bill Johnson is the materials manager at Bill’s Electric in Lexingoton, Kentucky. Bill is looking for ways to reduce inventories. A recent statement shows the following inventories: raw materials, $5,127,400; work-in-process, $6,432,000; and finished goods, $3,587,200. This year’s cost of goods sold will be about $34.7 million. Assume 52 business weeks per year. Answer following two questions: Show work

1A. What is the raw material inventory

1B. What is work in process inventory

1C. What is the finished goods inventory

1D. Average aggregate inventory value is

1E. Weekly sales is

1F. Express total inventory as weeks of supply.

1G. Express total inventory as Inventory turns.
Answer #1

Raw material inventory =$5,127,400

Work in process inventory =$6,432,000

Finished goods = $3,587,200

Average aggregate inventory value = Sum of above 3 = $5,127,400+$6,432,000+$3,587,200 =$15146600

Weekly sales = 34700000 / 52 = $667307.7

Weeks of Supply Total aggregate inventory / Weekly sales = 15146600 / $667307.7 = 22.70 weeks

Inventory turns = Cost of Sales / Total aggregate inventory =34700000 / 15146600 = 2.29 turns